OpenStudy (anonymous):

Find the next four terms in each arithmetic sequence. (2/3),(4/3),2,(8/3),.....

OpenStudy (anonymous):

\[\frac{4}{3} - \frac{2}{3} = \frac{2}{3} \] just keep adding 2/3 to yur sequence: 8/3+2/3 = 10/3 10/3+2/3 = 12/3 = 4 and so on..

OpenStudy (anonymous):

o ok thanks

OpenStudy (anonymous):

so after 12/3 it would be 14/3 and after that it would be 16/3

OpenStudy (anonymous):

yup, yu got it!